Calculate the amount of heat required to increase the temperature of 4 kg of water from 23 0C to 30 0C. (cw = 4180J/kg/ 0C)
1
Expert's answer
2020-12-03T10:58:35-0500

We can find the amount of heat required to increase the temperature of 4 kg of water from "23 ^{\circ}C" to "30 ^{\circ}C" from the formula:


"Q=mc\Delta T,"

"Q=4\ kg\cdot4180\ \dfrac{J}{kg\cdot\! ^{\circ}C}\cdot(30\ ^{\circ}C-23\ ^{\circ}C)=117040\ J."

Answer:

"Q=117040\ J."